# Hyperbolic Random Graphs: Clique Number and Degeneracy with Implications for Colouring

### v1

@@ 3 Degeneracy of Hyperbolic Random Graphs
A tool we make use of several times is the inner-ball of a point $x\in\mathcal{D}_{R}$, defined by $\mathcal{I}(x)=\mathcal{B}_{x}(R)\cap\mathcal{B}_{0}(r(x))$. Since area of an inner-ball is invariant under angular rotation we also write $\mathcal{I}(r)$. The inner-ball of a vertex is the inner-ball of the point using the vertex’ coordinates. The inner-neighbourhood of a vertex $u$ is the set of vertices (excluding $u$) contained in its inner-ball, that is, the neighbours of $u$ that have a smaller radius than $u$ (see Figure 2 a), and is denoted $\operatorname{\Gamma}(u)$.
We upper bound the degeneracy $\kappa(G)$ via the inner-neighbourhood by using the following informal process. Consider a graph $G$ and order its vertices $(v_{1},v_{2},\cdots,v_{n})$ by decreasing radius, so $r(v_{i})\geq r(v_{i+1})$, and iteratively remove vertices from $G$ one-by-one, from lower to higher index. Note that the set of neighbours of $v_{i}$ that have greater index than $i$ coincide with $\operatorname{\Gamma}(v_{i})$. This implies that the degree of each vertex $v_{i}$ at the time of its removal is $|\operatorname{\Gamma}(v_{k})|$. Let $u^{*}$ be the vertex $v_{k}$ that maximises $|\operatorname{\Gamma}(v_{k})|$. As the largest degree among vertices in this ordering is given by $|\operatorname{\Gamma}(u^{*})|$, we get the following upper bound for the degeneracy.
@@
We will now show that the largest inner-degree does not only yield an immediate upper bound on the degeneracy, but also a lower bound. Informally, this lower bound follows from the following argument. Order the vertices of the graph in ascending order of their radius, that is, $(v_{1},v_{2},\ldots,v_{n})$ such that $r(v_{i})\geq r(v_{i+1})$. For $i\in[n]$, let $G_{i}=G\setminus\{v_{1},v_{2},\ldots,v_{i}\}$. Let $v_{k}$ be the node with the maximum inner neighbourhood in $G$. We show (with a probabilistic guarantee) that graph $G_{k}$ has minimum degree $(1-o(1))|\operatorname{\Gamma}(v_{k})|$. Before we make this formal, we introduce a slightly modified version of [18, Lemma 3.3], that implies the following: the closer a vertex is to the origin, the more neighbours it has in expectation. This can be proven via the fact that the angle $\theta_{R}(r,x)$ is monotonically decreasing in $x$.
@@
From Corollary 3.1 it follows that any vertex with radius at most $r$ has in expectation at least as many neighbours with radius up to $r$, as the expected inner-degree of a vertex with radius exactly $r$. In order to derive a high-probability bound on $|\operatorname{\Gamma}(u^{*})|$ it now suffices to show concentration around the expectation of all considered neighbourhoods.
{observation}
Let $G\sim\mathcal{G}(n,\alpha,C)$ be a threshold HRG. Then, $\omega(G)\geq\sigma(G)$.
@@
For any subgraph $H\subseteq G$, it is clear that $\kappa(G)\geq\min_{v\in V(H)}\mathrm{deg}|_{H}(v)$, where $\mathrm{deg}|_{H}(v)=\sum_{w\in V\setminus\{v\}}\mathbb{I}\{w\leftrightarrow v,w\in V(H)\}$. We let $H$ be the (random) subgraph of $G$ created by restricting $G$ to vertices that land in $\mathcal{B}_{0}(r)$, that is, that have radius less or equal $r$, and keeping the same edges.
Then for any $v\in V$,
$\mathbb{E}[\mathrm{deg}|_{H}(v);r(v)\leq r]$ $=\sum_{w\in V\setminus\{v\}}\mathbb{P}(w\in\mathcal{B}_{0}(r)\cap\mathcal{B}_{r(v)}(R),r(v)\leq r)$
$\geq\sum_{w\in V\setminus\{v\}}\mathbb{P}(w\in\mathcal{B}_{0}(r)\cap\mathcal{B}_{r}(R),r(v)\leq r)$ (Corollary 3.1)
$=(n-1)\mu(\mathcal{B}_{0}(r)\cap\mathcal{B}_{r}(R))\mathbb{P}(r(v)\leq r).$
Thus $\mathbb{E}[\mathrm{deg}|_{H}(v)\,|\,r(v)\leq r]\geq(n-1)\mu(\mathcal{B}_{0}(r)\cap\mathcal{B}_{r}(R))\eqqcolon\gamma$. Since $\mathrm{deg}|_{H}(v)$ is a sum of Bernoulli random variables that are all independent under $\mathbb{P}(\,\cdot\,|\,r(v)\leq r)$, a Chernoff bound gives that it is highly concentrated around $\gamma$. Thus
$\mathbb{P}\Big{(}\min_{v\in V(H)}\mathrm{deg}|_{H}(v)<(1-o(1))\gamma\Big{)}$ $=\mathbb{P}\Big{(}\bigcup_{v\in V}\{\mathrm{deg}|_{H}(v)<(1-o(1))\gamma,r(v)\leq r\}\Big{)}$
$\leq n\mathbb{P}(\mathrm{deg}|_{H}(v)<(1-o(1))\gamma,r(v)\leq r)$
$\leq n\mathbb{P}(\mathrm{deg}|_{H}(v)<(1-o(1))\gamma\,|\,r(v)\leq r)\leq n\mathrm{e}^{-o(1)/2\gamma}.$
Taking $r$ to be the argmax of $\gamma$ yields $\gamma(r)\geq\mathbb{E}[\Gamma(u^{*})]$. By Section 3 and applying (2) at $R/2$ we have that $\gamma(r)\geq\mathbb{E}[\Gamma(u^{*})]\geq\mathbb{E}[\sigma(G)]\in n^{\Omega(1)}.$ Thus for any choice of $c$ there is a large enough $n$ such that
$\mathbb{P}(\kappa(G)<(1-o(1))|\operatorname{\Gamma}(u^{*})|)\leq\mathbb{P}(\kappa(G)<(1-o(1))\gamma(r))\leq n^{1-c},$
@@
In the rest of this section we derive bounds for the largest inner-degree on HRGs that hold w.e.h.p. which, by Sections 3 and 3.2, translate into results for the degeneracy. A vertex $v$ belongs to the inner-neighbourhood of a vertex $u$, if and only if it resides in the inner-ball of $u$. As such, we can use the measure $\mu\left(\mathcal{I}(r)\right)$ of an inner-ball $\mathcal{I}_{u}$ to bound the maximum inner-degree of a graph. Since the measure of the inner-ball is invariant under rotation around the origin, we write $\mu\left(\mathcal{I}(r)\right)$ instead of $\mu\left(\mathcal{I}_{u}\right)$ for $r=r(u)$. We sum up our results for the area of the inner-ball in the following lemma which is obtained by combining Lemmas C.1 and C.3 (see Appendix C) where explicit values for $\gamma$ and $\eta$ are also stated.
@@
$\mu(\mathcal{I}(r))=\left(1+\Theta\left(e^{-\alpha R}\right)\right)\frac{\alpha e^{-\alpha r}}{\alpha-1/2}\left(\gamma e^{\frac{1}{2}(2\alpha-1)(2r-R)}-\eta\right).$
@@
Theorem 3.5 (Bounds on degeneracy).
@@
Proof 3.6 (Proof sketch).
Differentiating $\mu(\mathcal{I}(r))$ we find $r^{*}$, which maximises it. The upper bound is obtained by using $r^{*}$ for the upper bound of the inner-ball with $(n-1)\mu(\mathcal{I}(r^{*}))$, and using a Chernoff bound along with a union bound which gives an upper bound for $|\operatorname{\Gamma}(u^{*})|$. Section 3 then yields the upper bound for the degeneracy. For the lower bound, we first show that there exists a vertex with a radius $\tilde{r}$ close in value to $r^{*}$. Then $(n-1)\mu(\mathcal{I}(\tilde{r}))$ lower bounds $|\operatorname{\Gamma}(u^{*})|$ w.e.h.p. This gives a lower bound for degeneracy, due to Lemma 3.2.
Recalling that $\omega(G)\leq\chi(G)\leq\kappa(G)+1$, and applying Sections 3 and 3.5, the following is immediate.
Corollary 3.7 (Bounds on chromatic number).
@@
Proof 3.9 (Proof sketch.).
Using a smallest-last vertex ordering [29] the number of colours required is upper-bounded by $\kappa(G)$. The linear running time is a direct consequence of HRGs being sparse. The approximation ratio is achieved by comparing the lower bound of the chromatic number in Corollary 3.7 to the upper bound of the degeneracy in Theorem 3.5.
@@ 4 Clique Number of Hyperbolic Random Graphs
For any graph $G$, the clique number $\omega(G)$, chromatic number $\chi(G)$, and degeneracy $\kappa(G)$ are related via the inequalities $\omega(G)\leq\chi(G)\leq\kappa(G)+1$. If $G$ has $\kappa(G)=\omega(G)+1$, then $\chi(G)$ has the same value and can be computed in linear time. For this reason we are interested in the relationship between clique number and degeneracy for hyperbolic random graphs. In Section 4.1 we show that the two differ and that for HRGs, the degeneracy is strictly larger than the clique number by a constant multiplicative constant. In Section 4.2 we give new insights about where in the hyperbolic disk the largest clique is formed. We then conclude the section by providing a new upper bound for the clique number that states a leading constant in front of the size of the core, and that is increasing in $\alpha$.
@@ 4.1 The gap between Clique Number and Degeneracy
Because of the centralising effect of hyperbolic geometry, one might hope to show that the clique contained in the core of the disk is the largest, and that $\omega(G)=(1-o(1))\kappa(G)$. This would achieve two things: it would also imply a tight bound for the chromatic number $\chi(G)$, sandwiching it between clique number and degeneracy. It would also imply a linear time $(1+o(1))$-approximation algorithm for both clique number and chromatic number.
In this section we disprove these claims. We show that there exists a constant gap between clique number and degeneracy; this is the content of Theorem 4.3. Before embarking on the details of the proof, we first sketch its idea. For any clique $K$, its vertex $u$ with largest radius, whose inner-degree $|\operatorname{\Gamma}(u)|$ bounds $|K|$. For $r(u)\in R/2+o(1)$ and $r(u)\in R/2+\omega(1)$, $|\operatorname{\Gamma}(u)|$ is already smaller by a multiplicative constant than the lower bound for the degeneracy given in Theorem 3.5. What remains is to extend the result to $r(u)\in R/2+\Theta(1)$, which requires more intricate arguments and is addressed in the following lemma.
@@
Let $G\sim\mathcal{G}(n,\alpha,C)$ be a threshold HRG, let $\Delta\in\Theta(1)$ be constant and let $K$ be a clique where $u\in K$ is the vertex with largest radius $r_{K}=R/2+\Delta$. Then w.e.h.p. there exists a constant $\varepsilon\in(0,1)$ such that $\omega(G)<(1-\varepsilon)\kappa(G)$.
@@
Figure 3: Illustration of a separator. (a) The line $a_{\varphi}$ partitions disk $\mathcal{D}_{R}$ into two halfdisks $\mathcal{D}_{1}$ (blue) and $\mathcal{D}_{2}$ (pink). The hypercycle $\mathcal{H}_{\varphi}$ (hatched area) is defined by the line $a_{\varphi}$. Points in located in different halfdisks and outside the hatched area have distance at least $R$. (b) The separator (hatched area) separates the inner-neighbourhood (grey area) of a vertex $u$ of radius $r_{K}$ into three sub-areas $\mathcal{S}_{0}$, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. Any vertex located in $\mathcal{S}_{1}$ has no edge to a vertex in $\mathcal{S}_{2}$.
Let $G^{\prime}$ be the inner-neighbourhood of $u$, i.e., the induced subgraph $G^{\prime}=G[\Gamma(u_{K})]$. Then $K\subseteq V(G^{\prime})$, and thus $\omega(G^{\prime})\geq|K|$. We show that $\omega(G^{\prime})<(1-\varepsilon)|\operatorname{\Gamma}(u_{K})|$ w.e.h.p.
We accomplish this by proving that there exists a colouring for $G^{\prime}$ with $(1-\varepsilon)|V(G^{\prime})|=(1-\varepsilon)|\operatorname{\Gamma}(u_{K})|$ many colours. This implies an upper bound for the chromatic number $\chi(G^{\prime})$, which also serves as an upper bound for $\omega(G^{\prime})$. We do this by partitioning the inner-neighbourhood $\mathcal{I}\left(r_{K}\right)$ into three disjoint sub-regions $\mathcal{S}_{0}\cup\mathcal{S}_{1}\cup\mathcal{S}_{2}$ such that no vertex in $\mathcal{S}_{1}$ is adjacent to any vertex in $\mathcal{S}_{2}$. Thus, $\mathcal{S}_{0}$ separates $\mathcal{S}_{1}$ from $\mathcal{S}_{2}$ and we can colour the set of vertices $V\cap\mathcal{S}_{1}$ with the same colours as $V\cap\mathcal{S}_{2}$. Thus if $\min\{|V\cap\mathcal{S}_{1}|,|V\cap\mathcal{S}_{2}|\}\geq\varepsilon|\operatorname{\Gamma}(u)|$ w.e.h.p., our desired statement will be proven.
@@
Using a Chernoff bound (Theorem A.1) for both $|V\cap\mathcal{S}_{1}|$ and $|V\cap\mathcal{S}_{2}|$, we obtain that neither random variable is smaller than $(1-1/\log(n))\mathbb{E}\left[|V\cap\mathcal{S}_{1}|\right]\geq(1-1/\log(n))^{-1}\varepsilon\mathbb{E}\left[|\operatorname{\Gamma}(u)|\right]$ w.e.h.p. On the other hand, another application of a Chernoff bound reveals $|\operatorname{\Gamma}(u)|\leq(1+1/\log(n))\mathbb{E}\left[|\operatorname{\Gamma}(u)|\right]$ w.e.h.p., since for $r(u)\in R/2+\Theta(1)$ we have $\mathbb{E}\left[|\operatorname{\Gamma}(u)|\right]\in\Theta(n^{(1-\alpha)})$ using Lemma 3.4. A union bound then shows that w.e.h.p., $\min\{|V\cap\mathcal{S}_{1}|,|V\cap\mathcal{S}_{2}|\}\geq\frac{\varepsilon\mathbb{E}\left[|\operatorname{\Gamma}(u)|\right]}{1-1/\log(n)}\geq\varepsilon|\operatorname{\Gamma}(u)|.$ As argued above, a naïve colouring that colours the vertices of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ with the same set of colours yields the upper bound.
@@
Let $G\sim\mathcal{G}(n,\alpha,C)$ be a threshold HRG. Then w.e.h.p. there exists a constant $\varepsilon\in(0,1)$ such that $\omega(G)<(1-\varepsilon)\kappa(G)$.
@@
Let $K$ be the largest clique of $G$, and let $u$ be the vertex of $K$ with maximal radius $r_{K}\coloneqq r(u)$. Recall that $\kappa(G)>(1+\delta)e^{-\alpha C/2}n^{1-\alpha}$ w.e.h.p. by Theorem 3.5.
Case 1 [$r\in R/2+\omega(1)$]: Observe that $K\subseteq V\cap\mathcal{I}\left(r_{K}\right)$. Hence, $|K|\leq|V\cap\mathcal{I}\left(r_{K}\right)|$. By Lemma 3.4 we get that
$\mu\left(\mathcal{I}\left(r_{K}\right)\right)\leq\left(1+\Theta\left(e^{-\alpha R}\right)\right)\frac{\alpha e^{-\alpha r}}{\alpha-1/2}\left(\gamma e^{\frac{1}{2}(2\alpha-1)(2r-R)}-\eta\right)\in\Theta(1)n^{-\alpha}e^{-(1-\alpha)\omega(1)},$
since $\gamma$ and $\eta$ are both constants. Taking the expectation and using a Chernoff bound (Theorem A.1) and a union bound over all vertices we have $|K|\in o(n^{1-\alpha})$ w.e.h.p. This is smaller by a constant gap than the degeneracy $\kappa(G)\in\Omega(n^{1-\alpha})$.

### v2

@@ 3 Degeneracy of Hyperbolic Random Graphs
A tool we make use of several times is the inner-ball of a point $x\in\mathcal{D}_{R}$, defined by $\mathcal{I}_{x}=\mathcal{B}_{x}(R)\cap\mathcal{B}_{0}(r(x))$. The inner-ball of a vertex is the inner-ball of the point using the vertex’ coordinates. The inner-neighbourhood of a vertex $u$ is the set of vertices (excluding $u$) contained in its inner-ball, that is, the neighbours of $u$ that have a smaller radius than $u$ (see Figure 2 a), and is denoted $\operatorname{\Gamma}(u)$.
We upper bound the degeneracy $\kappa(G)$ via the inner-neighbourhood by using the following informal process. Consider a graph $G$ and order its vertices $(v_{1},v_{2},\cdots,v_{n})$ by decreasing radius, so $r(v_{i})\geq r(v_{i+1})$, and iteratively remove vertices from $G$ one-by-one, from lower to higher index. Note that the set of neighbours of $v_{i}$ that have greater index than $i$ coincide with $\operatorname{\Gamma}(v_{i})$. This implies that the degree of each vertex $v_{i}$ at the time of its removal is $|\operatorname{\Gamma}(v_{i})|$. Let $u^{*}$ be the vertex $v_{k}$ that maximises $|\operatorname{\Gamma}(v_{k})|$. As the largest degree of a vertex at the time of its removal is given by $|\operatorname{\Gamma}(u^{*})|$, we get the following upper bound for the degeneracy.
@@
We will now show that the largest inner-degree does not only yield an immediate upper bound on the degeneracy, but also a lower bound. Informally, this lower bound follows from the following argument. Order the vertices of the graph in descending order of their radius, that is, $(v_{1},v_{2},\ldots,v_{n})$ such that $r(v_{i})\geq r(v_{i+1})$. For $i\in[n]$, let $G_{i}=G\setminus\{v_{1},v_{2},\ldots,v_{i}\}$. Let $v_{k}$ be the node with the maximum inner neighbourhood in $G$. We show (with a probabilistic guarantee) that the graph $G_{k}$ has minimum degree $(1-o(1))|\operatorname{\Gamma}(v_{k})|$. Before we make this formal, we introduce a slightly modified version of [18, Lemma 3.3], that implies the following: the closer a vertex is to the origin, the more neighbours it has in expectation. This can be derived via the fact that the angle $\theta_{R}(r,x)$ is monotonically decreasing in $x$.
@@
Corollary 3.1 tells us that any vertex with radius at most $r$ has in expectation at least as many neighbours up to radius $r$, as the expected inner-degree of a vertex with radius exactly $r$. In order to derive a high-probability bound on $|\operatorname{\Gamma}(u^{*})|$ it now suffices to show concentration around the expectation of all considered neighbourhoods.
@@
For any subgraph $H\subseteq G$, it is clear that $\kappa(G)\geq\min_{v\in V(H)}\mathrm{deg}|_{H}(v)$, where $\mathrm{deg}|_{H}(v)=\sum_{w\in V\setminus\{v\}}\operatorname{\mathds{1}}\{\{v,w\}\in E(H)\}$. We let $H$ be the (random) subgraph of $G$ created by restricting $G$ to vertices that land in $\mathcal{B}_{0}(r)$, that is, that have radius at most $r$, and keeping the same edges. Then for any $v\in V$,
$\mathbb{E}[\mathrm{deg}|_{H}(v)|r(v)]$ $=\sum_{w\in V\setminus\{v\}}\mathbb{P}(w\in\mathcal{B}_{0}(r)\cap\mathcal{B}_{r(v)}(R)\;|\;r(v))\operatorname{\mathds{1}}\{r(v)\leq r\}$
$=(n-1)\mu(\mathcal{B}_{0}(r)\cap\mathcal{B}_{r(v)}(R))\operatorname{\mathds{1}}\{r(v)\leq r\}$
$\geq(n-1)\mu(\mathcal{B}_{0}(r)\cap\mathcal{B}_{r}(R))\operatorname{\mathds{1}}\{r(v)\leq r\}.$ (Corollary 3.1)
We write $\gamma\coloneqq(n-1)\mu(\mathcal{B}_{0}(r)\cap\mathcal{B}_{r}(R))$. Since $\mathrm{deg}|_{H}(v)$ is a sum of Bernoulli random variables that are all independent under $\mathbb{P}(\,\cdot\,|\,r(v))$, a Chernoff bound (Theorem 2.1) gives
$\mathbb{P}\Big{(}\min_{v\in V(H)}\mathrm{deg}|_{H}(v)$ $<(1-\varepsilon)\gamma\Big{)}=\mathbb{P}\Big{(}\bigcup_{v\in V}\{\mathrm{deg}|_{H}(v)<(1-\varepsilon)\gamma,r(v)\leq r\}\Big{)}$
$\leq n\mathbb{P}(\mathrm{deg}|_{H}(v)<(1-\varepsilon)\gamma,r(v)\leq r)$ (Union bound)
$=n\mathbb{E}\left[\mathbb{P}(\mathrm{deg}|_{H}(v)<(1-\varepsilon)\gamma\;|\;r(v))\operatorname{\mathds{1}}\{r(v)\leq r\}\right]$ (Tower rule)
$\leq n\mathbb{E}\Big{[}\mathrm{e}^{-\varepsilon^{2}\mathbb{E}[\mathrm{deg}|_{H}(v)|r(v)]/2}\operatorname{\mathds{1}}\{r(v)\leq r\}\Big{]}$ (Chernoff bound)
$\leq n\mathrm{e}^{-\gamma\varepsilon^{2}/2}\mathbb{P}(r(v)\leq r)\leq n\mathrm{e}^{-\gamma\varepsilon^{2}/2}.$
Taking $r$ to be the argmax of $\gamma$ yields $\gamma(r)\geq\mathbb{E}[\Gamma(u^{*})]$. By Section 2 and applying (2) at $R/2$ we have that $\gamma(r)\geq\mathbb{E}[\Gamma(u^{*})]\geq\mathbb{E}[\sigma(G)]\in n^{\Omega(1)}.$ Thus, choosing $\varepsilon=1/\log(n)$, we obtain for any constant $c$ as long as $n$ is large enough
$\mathbb{P}(\kappa(G)<(1-\varepsilon)|\operatorname{\Gamma}(u^{*})|)\leq\mathbb{P}(\kappa(G)<(1-\varepsilon)\gamma(r))\leq n\mathrm{e}^{-\gamma(r)\varepsilon^{2}/2}\leq n\mathrm{e}^{-n^{\Omega(1)}/\log(n)}\leq n^{1-c},$
@@
In the rest of this section we derive bounds for the largest inner-degree on HRGs that hold w.e.h.p. which, by Sections 3 and 3.2, translate into results for the degeneracy. Since a vertex $v$ belongs to the inner-neighbourhood of a vertex $u$, if and only if it resides in the inner-ball of $u$, we can use the measure of an inner-ball $\mathcal{I}_{u}$ to bound the maximum inner-degree of a graph. Moreover, the measure of the inner-ball is invariant under rotation around the origin, which is why we write $\mu\left(\mathcal{I}(r)\right)$ instead of $\mu\left(\mathcal{I}_{u}\right)$ for $r=r(u)$. We sum up our results for the area of the inner-ball in the following technical lemma.
@@
$\mu(\mathcal{I}(r))$ $\geq\left(1+\Theta\left(e^{-\alpha R}\right)\right)\frac{\alpha e^{-\alpha r}}{(\alpha-1/2)}\left(\frac{2}{\pi}e^{\frac{1}{2}(2\alpha-1)(2r-R)}-\left(\frac{2}{\pi}-\frac{(\alpha-1/2)}{\alpha}\right)\right)$ and
where
Lemma 3.5.
Let $r^{*}$ be the radial coordinate of the point in $\mathcal{D}_{R}$ that maximises the measure of the inner-ball. Then $r^{*}=R/2+\log\left(\frac{\alpha\eta}{\gamma(1-\alpha)}\right)/(2\alpha-1)$.
Proof 3.6.
Using the expression of $\mu(\mathcal{I}(r))$ in Lemma 3.4,
@@
Theorem 3.7 (Bounds on the degeneracy).
@@
Proof 3.8 (Proof sketch).
The upper bound is obtained by using $r^{*}$ as stated in Lemma 3.5 for the upper bound of the inner-ball with $(n-1)\mu(\mathcal{I}(r^{*}))$, and using a Chernoff bound along with a union bound which gives an upper bound for $|\operatorname{\Gamma}(u^{*})|$ w.e.h.p. Section 3 then yields the upper bound for the degeneracy. For the lower bound, we first show that there exists a vertex with a radius $\tilde{r}$ close in value to $r^{*}$ w.e.h.p. Then $(1-o(1))(n-1)\mu(\mathcal{I}(\tilde{r}))$ lower bounds $|\operatorname{\Gamma}(u^{*})|$ w.e.h.p. This gives a lower bound for the degeneracy, due to Lemma 3.2.
Applying Sections 2 and 3.7, the following is immediate.
Corollary 3.9 (Bounds on the chromatic number).
@@
Proof 3.11 (Proof sketch.).
Using a smallest-last vertex ordering [30] the number of colours required is upper-bounded by $\kappa(G)$. Computing the smallest-last vertex ordering, and then using the ordering to colour the graph, both takes linear as the giant component is sparse w.e.h.p. by [24, Corollary 17]. The approximation ratio is achieved by comparing the lower bound of the chromatic number in Corollary 3.9 to the upper bound of the degeneracy in Theorem 3.7.
@@ 4 Clique Number of Hyperbolic Random Graphs
Recall that for any graph $G$, the clique number $\omega(G)$, chromatic number $\chi(G)$, and degeneracy $\kappa(G)$ are related via the inequalities $\omega(G)\leq\chi(G)\leq\kappa(G)+1$. For this reason we are interested in the relationship between clique number and degeneracy for hyperbolic random graphs. In Section 4.1 we show that the two differ and that for HRGs, the degeneracy is strictly larger than the clique number by a constant multiplicative constant. In Section 4.2 we give new insights about where in the hyperbolic disk the largest clique is formed. We then conclude the section by providing a new upper bound for the clique number in Section 4.3 that states a leading constant in front of the size of the core, and that is increasing in $\alpha$.
@@ 4.1 The gap between Clique Number and Degeneracy
Because of the centralising effect of hyperbolic geometry, one might hope to show that the clique contained in the core of the disk is the largest, and that $\omega(G)=(1-o(1))\kappa(G)$. This would achieve two things on HRGs: first, it would imply a tight bound for the chromatic number $\chi(G)$, sandwiching it between clique number and degeneracy. Moreover, it would also imply a linear time $(1+o(1))$-approximation algorithm for the two NP-complete problems clique number and chromatic number using a smallest-last vertex ordering (see [30]).
In this section we disprove these claims. We show that there exists a constant gap between clique number and degeneracy; this is the content of Theorem 4.3. Before embarking on the details of the proof, we first sketch its idea. For any clique $K$, its size is bounded by the inner-degree $|\operatorname{\Gamma}(u)|$, where $u$ is the vertex with largest radius among the vertices of $K$. For $r(u)\leq R/2+o(1)$ and $r(u)\in R/2+\omega(1)$, $|\operatorname{\Gamma}(u)|$ is already smaller by a multiplicative constant than the lower bound for the degeneracy given in Theorem 3.7. What remains is to extend the result to $r(u)\in R/2+\Theta(1)$, which requires more intricate arguments and is addressed in the following lemma.
@@
Let $G\sim\mathcal{G}(n,\alpha,C)$ be a threshold HRG, let $\Delta\in\Theta(1)$ and let $K$ be any clique where $u\in K$ is the vertex with largest radius $r_{K}=R/2+\Delta$. Then w.e.h.p. there exists a constant $\varepsilon\in(0,1)$ such that $|K|\leq(1-\varepsilon)|\operatorname{\Gamma}(u)|$.
@@
Figure 3: Illustration of a separator. (a) The line $a_{\varphi}$ partitions disk $\mathcal{D}_{R}$ into two halfdisks $\mathcal{D}_{1}$ (blue) and $\mathcal{D}_{2}$ (pink). The hypercycle $\mathcal{H}_{\varphi}$ (hatched area) is defined by the line $a_{\varphi}$. Points located in different halfdisks and outside the hatched area have distance at least $R$. (b) The separator (hatched area) separates the inner-neighbourhood (grey area) of a vertex $u$ of radius $r_{K}$ into three sub-areas $\mathcal{S}_{0}$, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. Any vertex located in $\mathcal{S}_{1}$ has no edge to a vertex in $\mathcal{S}_{2}$.
Let $G^{\prime}$ be the inner-neighbourhood of $u$, i.e., the induced subgraph $G^{\prime}=G[\Gamma(u)]$. Then $K\subseteq V(G^{\prime})$, and thus $\omega(G^{\prime})\geq|K|$. We show that $\omega(G^{\prime})<(1-\varepsilon)|\operatorname{\Gamma}(u)|$ w.e.h.p.
We accomplish this by proving that there exists a colouring for $G^{\prime}$ with $(1-\varepsilon)|V(G^{\prime})|=(1-\varepsilon)|\operatorname{\Gamma}(u)|$ many colours. This implies an upper bound for the chromatic number $\chi(G^{\prime})$, which also serves as an upper bound for $\omega(G^{\prime})$. We do this by partitioning the inner-neighbourhood $\mathcal{I}_{u}$ into three disjoint sub-regions $\mathcal{S}_{0}\cup\mathcal{S}_{1}\cup\mathcal{S}_{2}$ such that no vertex in $\mathcal{S}_{1}$ is adjacent to any vertex in $\mathcal{S}_{2}$. Thus, $\mathcal{S}_{0}$ separates $\mathcal{S}_{1}$ from $\mathcal{S}_{2}$ and we can colour the set of vertices $V\cap\mathcal{S}_{1}$ with the same colours as $V\cap\mathcal{S}_{2}$. Thus if $\min\{|V\cap\mathcal{S}_{1}|,|V\cap\mathcal{S}_{2}|\}\geq\varepsilon|\operatorname{\Gamma}(u)|$ w.e.h.p., our desired statement will be proven.
@@
Using a Chernoff bound for both $|V\cap\mathcal{S}_{1}|$ and $|V\cap\mathcal{S}_{2}|$, we obtain that neither random variable is smaller than $(1-1/\log(n))\mathbb{E}\left[|V\cap\mathcal{S}_{1}|\right]\geq(1-1/\log(n))^{-1}\varepsilon\mathbb{E}\left[|\operatorname{\Gamma}(u)|\right]$ w.e.h.p. On the other hand, another application of a Chernoff bound reveals $|\operatorname{\Gamma}(u)|\leq(1+1/\log(n))\mathbb{E}\left[|\operatorname{\Gamma}(u)|\right]$ w.e.h.p., since for $r(u)\in R/2+\Theta(1)$ we have $\mathbb{E}\left[|\operatorname{\Gamma}(u)|\right]\in\Theta(n^{(1-\alpha)})$ using Lemma 3.4. A union bound then shows that w.e.h.p., $\min\{|V\cap\mathcal{S}_{1}|,|V\cap\mathcal{S}_{2}|\}\geq\frac{\varepsilon\mathbb{E}\left[|\operatorname{\Gamma}(u)|\right]}{1-1/\log(n)}\geq\varepsilon|\operatorname{\Gamma}(u)|.$ As argued above, a naïve colouring that colours the vertices of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ with the same set of colours yields the upper bound.
@@
Let $G\sim\mathcal{G}(n,\alpha,C)$ be a threshold HRG. Then w.e.h.p. there exists a constant $\varepsilon\in(0,1)$ such that $\kappa(G)/\omega(G)>1+\varepsilon$.
@@
Let $K$ be the largest clique of $G$, and let $u$ be the vertex of $K$ with maximal radius $r_{K}\coloneqq r(u)$. We assume the following cases for $r_{K}$ which cover all possibilities:
Case 1 [$r_{K}\in R/2+\omega(1)$]: Observe that $K\subseteq V\cap\mathcal{I}\left(r_{K}\right)$. Hence, $|K|\leq|V\cap\mathcal{I}\left(r_{K}\right)|$ and
$\mu\left(\mathcal{I}\left(r_{K}\right)\right)\leq\left(1+\Theta\left(e^{-\alpha R}\right)\right)\frac{\alpha e^{-\alpha r_{K}}}{\alpha-1/2}\left(\gamma e^{\frac{1}{2}(2\alpha-1)(2r_{K}-R)}-\eta\right)\in\Theta(1)n^{-\alpha}e^{-(1-\alpha)\omega(1)},$
by Lemma 3.4 since $\gamma$ and $\eta$ are both constants. Taking the expectation and using a Chernoff bound we have $|K|\in o(n^{1-\alpha})$ w.e.h.p. Recall that $\kappa(G)>(1+\delta)e^{-\alpha C/2}n^{1-\alpha}$ w.e.h.p. by Theorem 3.7. It follows $\kappa(G)/|K|\in\omega(1)$ w.e.h.p.
Case 2 [$r_{K}\leq R/2+o(1)$]: Observe that the size of any clique $K\subseteq V\cap\mathcal{B}_{0}(r)$ is upper bounded by $X=|V\cap\mathcal{B}_{0}(r)|$. By Equation 2 we have $\mu(\mathcal{B}_{0}(r_{K}))\leq(1+o(1))n^{-\alpha}e^{-\alpha C/2}$. Hence, we get $\mathbb{E}[X]\leq(1+o(1))n^{1-\alpha}e^{-\alpha C/2}$. Since $X$ is a binomial random variable we can apply a Chernoff bound, which yields $|K|\leq X\leq(1+o(1))n^{1-\alpha}e^{-\alpha C/2}$ w.e.h.p. Taking $\varepsilon=\delta/(1+\delta)$ with $\delta$ as in Theorem 3.7, we have